OpenStudy (anonymous):

simplify: 10a^2b^3+30ab^2-20ab^3 / 5ab^2

OpenStudy (anonymous):

(10^a^2b^3 + 30ab^2 - 20ab^3) / 5ab^2 = (10ab^2 (ab + 3 - 2b) ) / 5ab^2 = 2(ab + 3 -2b) = 2{b(a-2) + 3}